The speed of a sparrow is x mph in still air. When the wind is blowing at 1 mph, the sparrow takes 5 hours to fly 12 miles to its nest and 12 hours to return again. He flies out directly into the wind and returns with the wind behind him. FInd his speed in still air.
Answer by mananth(16946)

The speed of a sparrow is x mph in still air.
When the wind is blowing at 1 mph,
speed will be x-1 mph against the wind
time taken 12 hours
..
Distance traveled both ways is the same
Distance = rate*time
12*(x-1) = 12
with wind
5*(x+1)=12
..
12(x-1)=5*(x+1)
12x-12=5x+5
12x-5x=12+5
7x=17
x= 17/7 mph in still air

the sparrow takes 5 hours to fly 12 miles to its nest and 12 hours to return again. He flies out directly into the wind and returns with the wind behind him. FInd his speed in still air.